# Uncertainty in latent representations of variational autoencoders optimized for visual tasks

### v2

@@ Abstract
Deep Generative Models (DGMs) can learn flexible latent variable representations of images while avoiding intractable computations, common in Bayesian inference. However, investigating the properties of inference in Variational Autoencoders (VAEs), a major class of DGMs, reveals severe problems in their uncertainty representations. Here we draw inspiration from classical computer vision to introduce an inductive bias into the VAE by incorporating a global explaining-away latent variable, which remedies defective inference in VAEs. Unlike standard VAEs, the Explaing-Away VAE (EA-VAE) provides uncertainty estimates that align with normative requirements across a wide spectrum of perceptual tasks, including image corruption, interpolation, and out-of-distribution detection. We find that restored inference capabilities are delivered by developing a motif in the inference network (the encoder) which is widespread in biological neural networks: divisive normalization. Our results establish EA-VAEs as reliable tools to perform inference under deep generative models with appropriate estimates of uncertainty.
@@
Keywords Uncertainty  $\cdot$ Variational Autoencoders  $\cdot$ Probabilistic Inference  $\cdot$ Deep Generative Models  $\cdot$ Divisive Normalization
Introduction
Variational autoencoders (VAEs) have been established as immensely valuable tools to learn about the statistics of data in an unsupervised manner and to interpret it through inferring latent variables [1, 2, 3, 4]. Attractively, latent variables of a VAE, $\mathbf{z}$, can be considered as generative factors underlying data $\mathcal{P}(\mathbf{x})$, which directly links inferences in the model to the interpretation of observations. Intuitively, this generative process is summarized as a generative model $\mathcal{P}(\mathbf{x}|\mathbf{z})$, which is complete with a prior distribution over latent features, $\mathcal{P}(\mathbf{z})$. VAEs thus establish a highly principled approach to inference, approximating the true Bayesian posterior distributions $\mathcal{P}(\mathbf{z}|\mathbf{x})$ by a recognition model $\mathcal{Q}(\mathbf{z}|\mathbf{x})$. Notably, VAEs simultaneously estimate the most likely values of latent features and the uncertainty associated with these values. Indeed, an adequate representation of uncertainty is key in multiple scenarios, including decision making, and sensory integration [5], which rely on knowledge about the uncertainty of the sources [6]. Critically for the healthcare domain, where diagnostics involve integrating results from multiple tests and imaging modalities, an adequate estimate of uncertainty is required [7]. Consequently, capitalizing on the promise of VAEs requires reliable estimates of uncertainty.
Focusing on visual inference tasks, we identify four fundamental requirements towards principled inference using a salient variable of natural image statistics: contrast [8]. First, at zero contrast, which carries no information, the inferred posterior needs to be equal to the prior (Figure 1a). Second, with increasing contrast, observations become progressively more informative and the posteriors should become increasingly distinct from the prior. Accordingly, the average distance between the means of the posteriors and the prior, which we term signal mean, should increase (Figure 1a and b, top). Third, more informative observations yield increasing variance in posteriors means (Figure 1a and b, center), usually called signal variance. Finally, with increasing contrast, more informative observations denote decreasing uncertainty about the latent generative component, i.e. decreasing posterior variance, referred to as noise variance (Figure 1a and b, bottom). However, when training a VAE on natural images (Figure 1c), although contrast-dependence of the posterior mean matches the requirements, contrast-dependent changes in posterior variances are at odds with the promise of a faithful representation of uncertainty (Figure 1b).
@@ Introduction
Fig. 1: Characteristic properties of inferred posteriors as contrast is varied. a, Expected behaviour of the posteriors of a pair of latent variables in natural images when contrast is systematically increased. Three example images (green, purple and mustard frames) are presented at varying contrast levels. As contrast increases, the inferred posteriors for each image (colored solid lines) progressively deviate both from the prior (dashed line) and from one another. There are two different sources of variance in latent representations. The signal variance refers to how the posterior mean changes with the input. The noise variance (or posterior variance) refers to the remaining uncertainty after having observed a given stimulus. Similarly, we call signal mean to the average distance between the prior’s mean and the posteriors’ mean. The noise mean will be zero in our models, as is usually assumed for VAEs, without loss of generality. b, Cartoon of expected behaviour of signal mean, signal variance and noise variance as a function of contrast showing qualitative trends of these quantities. c,d, Signal mean, signal variance and noise variance of the inferred latent posteriors in natural image-trained VAE (c) and EA-VAE (d, respectively), as a function of image contrast (Methods). Prior mean and variance are shown in dotted black lines. e, Comparison of the inference and reconstruction process in the VAE (left) and EA-VAE (right). In the VAE a single pool of latent variables $\mathbf{z}$ is inferred, while in the EA-VAE there is an additional global latent variable $s$ which acts multiplicatively on the latents, $\mathbf{z}$. f, Example test image patches and their respective reconstruction through VAE and EA-VAE. g, Receptive filters of example latents in the VAE and EA-VAE. h, Inferred posterior mean of the scaling variable for individual patches (dots) in the EA-VAE model, as a function of the measured contrast of these images.
@@
We argue that this signature of inaccurate inference is a consequence of the mismatch between the assumptions about the statistics of latent representations and those characteristic of generative factors underlying naturally occurring inputs. In its basic formulation, latent variables in VAEs are assumed to be either independent or linearly correlated. Natural stimuli, however, have been shown to display strong nonlinear dependencies between latents [9]. Thus, we propose a theoretically well-motivated and powerful update to the standard VAE architecture, the Explaining-Away-VAE (EA-VAE for short). EA-VAE incorporates a global multiplicative scaling variable $s$ in the generative model, which is inferred by the recognition model (Figure 1e) simultaneously with the traditional latent variables $\mathbf{z}$. This extension was inspired by the Gaussian Scale Mixtures (GSM) model [9], a simple yet powerful model of natural images for image compression and denoising, which has also been successful in interpreting behavioral and neural data in biological perception [10, 11, 12, 13]. The introduced scaling variable can replace correlated modulation of latent activations upon changes in image contrast, as recruiting the scaling variable can explain global changes in latent activations. Hence the term explaining-away in our EA-VAEs. In contrast to a standard VAE model, the EA-VAE produces contrast-dependent posteriors on natural images that match the fundamental requirements of probabilistic inference (Figure 1d).

### v3

# Remedying uncertainty representations in visual inference through Explaining-Away Variational Autoencoders

@@ Abstract
Optimal computations under uncertainty require an adequate probabilistic representation about beliefs. Deep generative models, and specifically Variational Autoencoders (VAEs), have the potential to meet this demand by building latent representations that learn to associate uncertainties with inferences while avoiding their characteristic intractable computations. Yet, we show that it is precisely uncertainty representation that suffers from inconsistencies under an array of relevant computer vision conditions: contrast-dependent computations, image corruption, out-of-distribution detection. Drawing inspiration from classical computer vision, we present a principled extension to the standard VAE by introducing a simple yet powerful inductive bias through a global scaling latent variable, which we call the Explaining-Away VAE (EA-VAE). By applying EA-VAEs to a spectrum of computer vision domains and a variety of datasets, spanning standard NIST datasets to rich medical and natural image sets, we show the EA-VAE restores normative requirements for uncertainty. Furthermore, we provide an analytical underpinning of the contribution of the introduced scaling latent to contrast-related and out-of-distribution related modulations of uncertainty, demonstrating that this mild inductive bias has stark benefits in a broad set of problems. Moreover, we find that EA-VAEs recruit divisive normalization, a motif widespread in biological neural networks, to remedy defective inference. Our results demonstrate that an easily implemented, still powerful update to the VAE architecture can remedy defective inference of uncertainty in probabilistic computations.
@@
Keywords Uncertainty $\cdot$ Deep generative models $\cdot$ NeuroAI $\cdot$ Probabilistic Inference $\cdot$ Explaining-Away $\cdot$ Divisive Normalization
1 Introduction
An adequate representation of uncertainty is key to reliable decision making and sensory integration [28], and thus its importance for artificial systems is increasingly evident as flexibility, adaptation and multimodality become widespread requirements. Maintaining an estimate of the uncertainty associated with inferences permits better reasoning about beliefs [21], generalization [12], and assessment of risks associated with alternative decisions [1]. Furthermore, it allows for more efficient use of acquired knowledge for learning [11, 32, 20], and downstream computations in a hierarchical processing pipeline, among others. In particular, uncertainty-aware AI systems are argued to be necessary for exploring novel domains [45, 22]. Indeed, probabilistic machine learning promises to deliver tools that are not only capable of inferring best estimates but also the associated confidence in those estimates [14]. In this context, generative models have been advocated as a central framework for probabilistic machine learning [3, 41]. Recently, deep learning has been elevated to accommodate deep generative models, including generative adversarial networks [16], variational autoencoders (VAEs) [24, 33], and diffusion models [23, 40, 39]. In particular, when learning a latent representation of the data, VAEs provide an explicit representation of the uncertainty of inferences, rendering them a central tool for applications that require probabilistic computations [24, 18, 27, 42].
@@ 1 Introduction
Fig. 1: Characteristic properties of inferred posteriors as contrast is varied. a, Expected behavior of the posteriors of a pair of latent variables in natural images when contrast is systematically increased. Three example images (green, purple and mustard frames) are presented at varying contrast levels. As contrast increases, the inferred posteriors for each image (colored solid lines) progressively deviate both from the prior (dashed line) and from one another. There are two different sources of variance in latent representations. The signal variance refers to how the posterior mean changes with the input. The noise variance (or posterior variance) refers to the remaining uncertainty after having observed a given stimulus. Similarly, we call signal mean the average distance between the prior’s mean and the posteriors’ mean. The noise mean will be zero in our models, as is usually assumed for VAEs, without loss of generality. b, Cartoon of expected behavior of signal mean, signal variance and noise variance as a function of contrast showing qualitative trends of these quantities. c,d, Signal mean, signal variance and noise variance of the inferred latent posteriors in natural image-trained VAE (c) and EA-VAE (d, respectively), as a function of image contrast (Supplementary Section˜2.2.2). Prior mean and variance are shown in horizontal dashed black lines, and observation noise is shown in vertical dotted black lines. e, Comparison of the inference and reconstruction process in the VAE (left) and EA-VAE (right). In the VAE a single pool of latent variables $\mathbf{z}$ is inferred, while in the EA-VAE there is an additional global latent variable $s$ which acts multiplicatively on the latents, $\mathbf{z}$. f, Example test image patches and their respective reconstruction through VAE and EA-VAE. g, Receptive filters of example latents in the VAE and EA-VAE. High-saturation boxes surround informative latent units while low saturation boxes surround non-informative ones (Supplementary Section˜2.2.1). h, Inferred posterior mean of the scaling variable for individual patches (dots) in the EA-VAE model, as a function of the measured contrast of these images.
A key element of probabilistic computations is that evidence, the likelihood, is combined with expectations, the prior, to interpret observations, yielding the posterior distribution whose width carries information about uncertainty of inferences. Given the central role of uncertainty in VAEs, it is crucial to evaluate to what extent uncertainty representations fulfill elementary requirements set by probability theory. Indeed, the capacity of deep generative models to faithfully represent uncertainty has been questioned [29]. Focusing on visual inference tasks, we first consider contrast, a salient variable of natural image statistics, and identify four fundamental requirements towards principled inference [2]. First, at zero contrast, where the image carries no information, the inferred posterior needs to be equal to the prior (Figure˜1a).
Second, with increasing contrast, observations become progressively more informative and the corresponding posteriors should become increasingly distinct from the prior. Accordingly, the average distance between the means of the posteriors and the prior, which we term signal mean, should increase (Figure˜1a and b, top). Third, more informative observations yield increasing variance in posteriors means (Figure˜1a and b, center), usually called signal variance. Finally, with increasing contrast observations become more informative, resulting in decreasing uncertainty about the latent generative component, i.e. decreasing posterior variance, referred to as noise variance (Figure˜1a and b, bottom). However, in this work we show that when training a VAE on natural images (Figure˜1f), although contrast-dependence of the posterior mean matches these requirements, contrast-dependent changes in posterior variances are at odds with the promise of a faithful representation of uncertainty (Figure˜1, cf. b and c).
@@
We argue that this signature of inaccurate inference is a consequence of the mismatch between the assumptions about the statistics of latent representations and those characteristic of generative factors underlying naturally occurring inputs. In its basic formulation, latent variables in VAEs are assumed to be either independent or linearly correlated. Natural stimuli, however, have been shown to display strong nonlinear dependencies between latents [44].
We propose a theoretically well-motivated and powerful update to the standard VAE architecture, the Explaining-Away-VAE (EA-VAE for short). EA-VAE incorporates an inductive bias in the generative model by learning an extra latent variable, which acts as a global multiplicative scaling that is capable of modulating the variance of the data. This extension was inspired by the Gaussian Scale Mixtures (GSM) model [44], a simple yet powerful model of natural images for compression and denoising, which has also been successful in interpreting behavioral and neural data in biological perception [37, 31, 10, 9]. As recruiting the scaling variable can explain global changes in the input, the introduced scaling variable can replace correlated modulations of latent activations upon changes in image contrast. Hence the term explaining-away in our EA-VAEs.